DQuestion 4 3 pts A stock has a price of $12.59 and earnings of $1. What is its price/earnings ratio? D Question 5 3 pts A firm is priced at $38.68, with a cost of capital of 14% and expected earnings of $1.63. What is its PVGO? D Question 6 3 pts What is the dividend-yield of this firm? Sales Dividends Market Value Earnings (NI) $122,000 $13,000 $208,000 $36,000Please answer them Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Answer:

Question 4)

Price earning ratio= Price/Earnings=12.59/1=12.59

Question 5)

PVGO=Stock price- earnings / Cost of capital

PVGO=38.68-1.63/14%=$27.0371

Question 6)

Dividend yield = Dividend /Market Value= 13000/208000=0.0625=6.25%